OpenStudy (anonymous):

A man can swim with a speed of 4.0 km/h in still water. How long does he take to cross a river 1.0 km wide if the river flows steadily at 3.0 km/h and he makes his strokes normal to the river current? How far down the river does he go when he reaches the other bank?

OpenStudy (anonymous):

the relative motion of the man with respect to the river is shown in the fig 

OpenStudy (anonymous):

so he moves with a relative speed of square root of (3^2+4^2)=5km/h then: t=d/v = 1/5 =.2 hours... after .2 hours he would move a horizontal distance of 3*.2 = .6 km that's what I think

OpenStudy (anonymous):

oh, but the answer is 750m and the time is 15 min

OpenStudy (anonymous):

pls let me know if u could get the answer as the above one

OpenStudy (anonymous):

huh? what I am thinking of right now is if it the time is really 15 mins, then the total speed is 4 km/h thus the motion is not relative! ,, and according to that the horizontal distance is d=3*1/4= 3/4 km= 750 meters. does still water mean the water that does not move?

OpenStudy (radar):

As previously shown the resultant speed is 5 km/h. Next compute the angle that this rate is from the normal to the river. Now compute distance x, Tan 36.87 = x, x=.75km or 750 meters.